# Low Fidelity Visuo-Tactile Pretraining Improves Vision-Only Manipulation Performance

### v3

@@ 3.3.3 Action Chunking Transformers
4 Experimental Results
@@
To examine the effect of pretraining with a low-fidelity tactile sensor on imitation learning agents, we tested our trained ACT and diffusion agents on the cable plugging task described in Section 3.2, following the same evaluation criteria used in [3] so that we can directly compare our results. For each evaluation, the agent attempted the task 20 times, with success being recorded if the agent plugged in and then released the USB. To increase the difficulty of the task, random noise was applied to the goal action, sampled from a normal distribution with a standard deviation of 2.5 mm.
@@
For both ACT and Diffusion Policy, we evaluated the effect of low fidelity visuo-tactile pretraining on both vision only agents and visuo-tactile agents. Our results, along with the evaluation using GelSight instead of BeadSight, can be seen in Table 1.
In the prior visuo-tactile pretraining experiments using GelSight, pretraining was found to increase the performance of both vision-only and visuo-tactile agents for both ACT and diffusion policy, although the improvement was more pronounced with the vision-only agents. However, replacing the GelSight with a low-cost, low-fidelity sensor reversed part of this trend. We found that pretraining significantly harmed the visuo-tactile agent’s performance, halving the success rate of both ACT and Diffusion Policy.
Pretrained,
Tactile
+ Vision
Pretrained,
Vision Only
Tactile
+ Vision
Vision Only
Pretrained,
Tactile**
+ Vision
Tactile**
+ Vision
Diffusion (B) 0.05 0.7 0.1 0.05 0.8 0.2
ACT (B) 0.25 0.65 0.5 0.15 - -
Diffusion (G) 0.75 0.75 0.7 0.45 - -
ACT (G) 0.95 0.85 0.9 0.2 - -
Table 1: Results of the USB cable plugging experiment, with (B) denoting the BeadSight and (G) showing the previous results with GelSight from [3]. ** Denotes that the tactile encoder is frozen.
@@
Figure 4: tSNE plot showing the distribution shift between training and deployment for the tactile observation.
In contrast, visuo-tactile pretraining significantly improved the performance of the vision-only agent, allowing the vision-only policies to outpace the visuo-tactile policies for both ACT and Diffusion agents. We believe this counter-intuitive result, that less information led to an improvement in model performance, is likely due to the network overfitting to the BeadSight observations. Not only is the BeadSight stochastic, but the beads can undergo significant drift between experiments, leading to a train-test distribution mismatch. Also, overfitting would explain why Diffusion Policy is more negatively affected by BeadSight observation than ACT, as the temporal ensembling in ACT helps address out-of-distribution states. This resistance is especially important in tasks like cable plugging, where a single wrong action (such as opening the gripper at the wrong time) can lead to an irrevocable failure.
To investigate BeadSight overfitting, we compared the BeadSight observations collected during training with those collected during model testing. We extracted the relevant tactile features using the tactile encoder from the pretraining step and graphed a tSNE plot [38] of the resulting BeadSight embeddings. This plot can be seen in Figure 4. The tSNE results show a clear divide between the training data set and the bead sight observations collected during testing, reflecting a significant train-test distribution shift, which can cause overfitting.
To address the agent’s overfitting to the BeadSight data, we froze the bead sight encoder weights before training the agents. For the pre-trained agent, these frozen weights were the result of the contrastive pretraining, while for the non-pretrained agent, the weights were frozen at initialization. Although this approach works for diffusion, where the encoder returns a single 512 vector, ACT removes the final pooling layer from the encoder passing the entire feature map (a 225x512 sequence) into the transformer, which we can’t freeze. To properly reduce overfitting via encoder freezing, we would need a narrower encoder bottleneck, which would require a significant alteration of the model. Therefore, to preserve consistency, we only evaluated this approach on the Diffusion model. The performance of these policies can be seen in Table 1. Our results show that freezing the weights reduced overfitting, causing a significant improvement in performance over the fine-tuned models and overtaking the performance of their respective vision-only and visuo-tactile counterparts.
@@
Comparing these final results to those achieved in [3] using GelSight, we see that without pretraining, using a low-quality tactile sensor for imitation learning leads to a significant drop in performance. However, visuo-tactile pretraining, when combined with weight freezing, can help close this gap. When a vision-only policy is desired, then an inexpensive tactile sensor can be used.
5 Conclusion
@@
In this work, we explored combining a low-fidelity tactile sensor (BeadSight) with pretraining for imitation learning. We found that using low-fidelity tactile sensors for imitation learning (without pretraining) results in a significant reduction in performance and is prone to overfitting. However, this performance deficit can be significantly reduced using contrastive pretraining. Also, by freezing the pre-trained weights, much of the overfitting issues can be resolved. Additionally, by using a low-fidelity tactile sensor and visuo-tactile pretraining, we were able to significantly improve the performance of a vision-only agent, replicating a prior finding that used a high-resolution GelSight tactile sensor.
@@
These relative performance improvements used $\sim$ 100 expert demonstrations and low computational intensity imitation learning methods. This suggests tactile pretraining can be another tool for improving manipulation performance on difficult tactile-rich tasks with efficient computational methods. Other work in IL has attempted to dramatically reduce the number of demonstrations necessary for an agent to imitate an action [39] and one can imagine tactile pretraining as a way to further improve IL’s efficiency on tactile-rich tasks.
@@
Future work could explore adding on to the imitation learning pipeline proposed here. In addition to freezing the tactile encoder after pretraining, we may also explore encoding the tactile image in a latent space to improve tactile performance at inference. A latent vector produced by an autoencoder could be used as the tactile observation rather than the image of the bead sac directly. Tracking and masking out the image of the tactile sensor could allow any gripper finger used in deployment, improving utility of the vision-only agent. Finally, the benefit of freezing the encoder together with the pretraining approach implies a way to pretrain an encoder on a large, diverse dataset and use this to inform downstream tactile manipulation tasks, with only vision used at execution. This kind of pretrained model at an appropriate scale could broadly improve general robot manipulation.
Acknowledgments

### v4

@@ III-C3 Action Chunking Transformers
IV EXPERIMENTAL RESULTS
IV-A Cable Plugging Task
TABLE I: Success rates: USB cable plugging experiment
Diffusion
Diffusion
ACT
Diffusion
ACT
Pre-trained,
Tactile + Vision
0.05 0.8 0.25 0.75 0.95
Pre-trained,
Vision
0.7 - 0.65 0.75 0.85
No Pre-training,
Tactile + Vision
0.1 0.2 0.5 0.7 0.9
No Pre-training,
Vision
0.05 - 0.15 0.45 0.2
@@ IV-A Cable Plugging Task
To examine the effect of pre-training with a low-fidelity tactile sensor on imitation learning agents, we tested our trained ACT and Diffusion agents on the cable plugging task described in Section III-B, following the same evaluation criteria used in [3] so that we can directly compare our results. For each evaluation, the agent attempted the task 20 times, with success being recorded if the agent plugged in and then released the USB. To increase the difficulty of the task, random noise was applied to the executed actions, sampled from a normal distribution with a standard deviation of 2.5 mm.
@@
For both ACT and Diffusion Policy, we evaluated the effect of low-fidelity visuo-tactile pre-training on both vision-only and visuo-tactile agents. Results are shown in Table I where “Gel” denotes the GelSight sensor and all other results use the BeadSight sensor. “Frozen” denotes freezing the tactile encoder (not updating the encoders weights) during downstream training.
Prior work using GelSight demonstrated that visuo-tactile pre-training improves the performance of both vision-only and visuo-tactile agents, with the most notable gains observed in the vision-only setting. When replacing GelSight with BeadSight, an open-source, low-fidelity tactile sensor, we observed a shift in this trend: while pre-training with BeadSight negatively impacted visuo-tactile agents, reducing the success rate of both ACT and Diffusion Policy, its effect on vision-only agents was highly beneficial. Specifically, vision-only policies trained with BeadSight pre-training outperformed their visuo-tactile counterparts, demonstrating that even a low-cost sensor can still provide valuable training signals.
The gap between visuo-tactile and vision-only agents likely stems from over-fitting to BeadSight’s stochastic and variable observations. Unlike GelSight, BeadSight’s beads can drift significantly between experiments, leading to a train-test distribution mismatch. This explains why visuo-tactile agents struggled, particularly with Diffusion Policy, which lacks ACT’s temporal ensembling that mitigates out-of-distribution issues. This robustness is crucial in high-stakes tasks like cable plugging, where a single incorrect action can lead to irreversible failure.
To investigate BeadSight over-fitting, we compared its training and testing observations using a tSNE plot [38], generated from tactile embeddings extracted by the pre-trained encoder (Figure I). The results confirm a clear distribution shift between training and testing, highlighting the challenge of direct visuo-tactile policy transfer with BeadSight.
@@
Figure 6: tSNE plot showing the distribution shift between training and deployment of the tactile observation.
To mitigate this, we froze the BeadSight encoder weights before training the agents. For pre-trained models, we froze the weights obtained from contrastive pre-training; for non-pretrained models, we froze them at initialization. This approach effectively reduced over-fitting in Diffusion Policy, leading to substantial performance gains, surpassing both fine-tuned models and their respective vision-only and visuo-tactile baselines. However, due to ACT’s reliance on full feature maps rather than a compressed embedding, freezing the encoder was infeasible without major architectural modifications. Therefore, we only evaluated weight freezing on Diffusion Policy.
@@
Comparing these results to those achieved in [3] using GelSight, we find that without pre-training, using a low-fidelity tactile sensor in imitation learning leads to a performance drop. However, when combined with visuo-tactile pre-training and weight freezing, the performance gap can be significantly reduced. Most importantly, when a vision-only policy is preferred, BeadSight offers a low-cost yet highly effective way to enhance learning, demonstrating that even open-source tactile sensors can provide meaningful benefits in policy training.
IV-B Drawer Task
TABLE II: Success rates: drawer experiment, removed tactile encoder
Opened Picked Placed Closed
No Pre-training 0.70 0.35 0.20 0.20
Vision Pre-training 0.45 0.25 0.05 0.05
Dissimilar Pre-training 0.90 0.70 0.50 .50
Similar Pre-training 1.00 0.80 0.50 .50
Identical Pre-training 1.00 0.90 0.55 0.55
The drawer task evaluation consisted of the long-horizon drawer pick-and-place task. Each agent attempted the task 20 times. Success was assessed across the different sub-tasks (opening drawer, picking block, placing block in drawer, closing drawer). Each sub-task is only marked successful if the previous task was successful, meaning that the success rate for closing the drawer is also the total success rate for the overall task completion. Diffusion Policy was chosen for this task, and the effect of freezing vs. ablating the tactile encoder was compared.
The supporting drawer task (similar) and original USB task (dissimilar) were used to investigate the effects of pre-training with different tasks. We also evaluated a vision pre-training ablation, which pre-trained without the tactile sensor to investigate whether the mere alignment of different camera views contributed to the performance improvement. For this ablation, all fixed camera images were aligned to the end-effector camera image and the end-effector position. This mimicked the visuo-tactile contrastive pre-training structure, where camera images were aligned to the tactile sensor images and end-effector position.
The best-performing agent for this task removed, rather than froze, the tactile encoder before downstream training, as shown in Table II. Both no pre-training and vision-only pre-training resulted in a 60–90% performance decline compared to visuo-tactile pre-training across identical, similar, and dissimilar tasks. This underscores the essential role of tactile information in providing critical feedback for manipulation, and highlights that purely vision agents struggle without the rich contact information that tactile sensing offers. Furthermore, visuo-tactile pre-training on any of these tasks yielded comparable performance improvements, suggesting that a pre-trained visuo-tactile encoder can enhance any vision-only agent downstream, assuming the same robot system morphology.
The frozen tactile encoder agent (Table III) produced mixed results, although the same technique led to an improvement for the previous cable plugging task. We speculate that the effect of freezing compared to ablating the tactile encoder depends on the drift of various physical parameters of the sensor, making this choice a hyper-parameter that can be tested to tune performance on different tasks. However, ablating the encoder is likely to be more stable, since as long as the sensor remains visually similar in the image, any drift on the tactile parameters will not affect the inference result.
TABLE III: Success rates: drawer experiment, frozen tactile encoder
Opened Picked Placed Closed
Dissimilar Pre-training 0.45 0.00 0.00 0.00
Similar Pre-training 0.95 0.60 0.50 0.45
Identical Pre-training 0.60 0.35 0.25 0.20
V CONCLUSION
@@
In this work, we explored combining a low-fidelity tactile sensor (BeadSight) with pre-training for imitation learning. By using a low-fidelity, open-source tactile sensor and visuo-tactile pre-training, we were able to significantly improve the performance of a vision-only agent, replicating a prior finding that used a high-resolution GelSight tactile sensor. We show that freezing the tactile encoder after pre-training can be used to mitigate train-test distribution shifts when using a low-fidelity sensor, and the choice between freezing or ablating the sensor can act as a hyper-parameter (although ablating is likely more stable).
Our follow-up investigation further supported this visuo-tactile pre-training technique. We showed that pre-training with vision alone could not lead to the same performance improvement and that visuo-tactile pre-training on a dissimilar task could benefit the downstream vision-only agent.
@@
These relative performance improvements used $\sim$ 100 expert demonstrations and low computational intensity imitation learning methods. This suggests tactile pre-training can be another tool for improving manipulation performance on difficult tactile-rich tasks with efficient computational methods. Other work in IL has attempted to dramatically reduce the number of demonstrations necessary for an agent to imitate an action [39, 25, 40] and one can imagine tactile pre-training as a way to further improve IL’s efficiency on tactile-rich tasks.
Future work could explore adding to the imitation learning pipeline proposed here. Currently, though the tactile sensor is not active during inference, it still needs to be physically attached to the end-effector to maintain consistent visual observations. Tracking and masking out the image of the tactile sensor could allow any gripper finger used in deployment, improving the utility of the vision-only agent. There is also the critical assumption of identical robot system morphology, which prevents learning across physically different robot platforms. This constraint could potentially be mitigated by exploring latent space pre-training approaches that leverage a larger visuo-tactile dataset, collected across a wide range of robot embodiments.
@@
Finally, the benefit of freezing the encoder together with the pre-training approach implies a way to pre-train an encoder on a large, diverse dataset and use this to inform downstream tactile manipulation tasks, with only vision used at execution. This kind of pretrained model at an appropriate scale could broadly improve general robot manipulation.
APPENDIX
We also include data from using a visuo-tactile agent with a fine-tuned tactile encoder for the drawer experiment in Table IV.
In the original USB plugging experiment in Table I, this was denoted “Tactile + Vision”, where the agent used both vision and tactile observations, and the tactile encoder was fine-tuned when training the imitation learning agent (not frozen). Although this approach was already shown to perform worse than the vision-only agent when using BeadSight, we collected this data for the drawer experiment and included it for completion. Note that [3] demonstrated that with a more precise commercial sensor such as GelSight, the visuo-tactile agent out-performs the vision-only agent on the USB cable plugging task.
TABLE IV: Success rates: drawer experiment, finetuned tactile encoder
Opened Picked Placed Closed
Dissimilar Pre-training Task 1.00 0.55 0.35 0.30
Similar Pre-training Task 0.85 0.45 0.10 0.10
Identical Pre-training Task 0.75 0.50 0.25 0.25